# Bayesian Estimation and Tuning-Free Rank Detection for Probability Mass Function Tensors

### v1

@@ III Variational Bayesian PMF Estimation
Estimates of the CPD components $\bm{\lambda}$ and $\{\bm{A}_{n}\}_{n=1}^{N}$ (and, therefore, of the joint PMF tensor $\bm{\mathcal{X}}$) can be readily obtained from $\bm{Y}$ by first obtaining empirical estimates of third-order marginal PMF tensors from $\bm{Y}$, followed by coupled tensor factorization [5], [7]. Alternatively, a maximum likelihood (ML) estimate of the PMF tensor can be obtained by fitting the observed data to the CPD model [11], [12]. While these approaches work quite well, they assume that the CPD rank $R$ is known, a scenario which rarely arises in practice. We therefore seek to formulate the problem within the Bayesian paradigm and design an algorithm to estimate the CPD components and automatically detect the rank of $\bm{\mathcal{X}}$.
@@ III-A Model Specification
The Bayesian model is specified by first identifying the model parameters than one wishes to estimate, finding an expression for the likelihood of the observations given the model parameters, and then assigning prior distributions to the model parameters.
@@
$p\big{(}\bm{Y}\,\big{|}\,\bm{\lambda},\{\bm{A}_{n}\}_{n=1}^{N}\big{)}=\prod_{t=1}^{T}\sum_{r=1}^{R}\lambda_{r}\prod_{n=1}^{N}\bm{A}_{n}(y_{n,t},r).$ (4)
The naïve Bayes model assumption allows us to introduce a latent variable $\bm{z}\in\mathbb{R}^{R}$ which contains the hidden structure that governs the observation $\bm{y}$. Let $\bm{z}=[z_{1},\dotsc,z_{R}]^{\mathsf{T}}$ where a particular element $z_{r}$ is equal to 1 and all other elements are equal to 0. Therefore, the elements of $\bm{z}$ satisfy $z_{r}\in\{0,1\}$ and $\sum_{r=1}^{R}z_{r}=1$. Each observation $\bm{y}$ is associated with a latent variable $\bm{z}$ which identifies the latent component from which $\bm{y}$ (and, therefore, $\bm{x}$) was drawn. In other words, if $\bm{y}$ is drawn from component $s\in\{1,\dotsc,R\}$ then $z_{r}=1$ if $s=r$ and $z_{r}=0$ if $s\neq r$. Defining $\bm{Z}=[\bm{z}_{1},\dotsc,\bm{z}_{T}]$ where $\bm{z}_{t}=[z_{1,t},\dotsc,z_{r,t}]^{\mathsf{T}}$, we can express the conditional distribution of $\bm{Y}$ given $\bm{Z}$ and $\{\bm{A}_{n}\}_{n=1}^{N}$ (up to a constant) as
@@
Next, we specify prior distributions for the component parameters $\bm{\lambda}$ and $\{\bm{A}_{n}\}_{n=1}^{N}$, which are subject to the probability simplex constraints (see Section II). An appropriate distribution for variables defined on a probability simplex is the Dirichlet distribution, which is a multivariate distribution over a set random variables that are subject to nonnegativity and sum-to-one constraints. In particular, given $K$ random variables $\bm{\mu}=[\mu_{1},\dotsc,\mu_{K}]^{\mathsf{T}}$ where $0\leq\mu_{k}\leq 1$, $k=1,\dotsc,K$ and $\sum_{k=1}^{K}\mu_{k}=1$, the Dirichlet distribution is given by (e.g., [1])
$\mathrm{Dir}(\bm{\mu}\,|\,\bm{\alpha})=C(\bm{\alpha})\prod_{k=1}^{K}\mu_{k}^{\alpha_{k}-1},~{}~{}\alpha_{k}>0{\color[rgb]{0,0,0},}~{}\forall k,$ (7)
where $C(\bm{\alpha})$ is the normalization constant, defined in terms of the standard gamma function [^2] For any real number $x$, the standard gamma function is defined as $\Gamma(x)=\int_{0}^{\infty}t^{x-1}\mathrm{e}^{-t}\mathrm{d}t$. $\Gamma(\cdot)$ as
@@
and $\bm{\alpha}=[\alpha_{1},\dotsc,\alpha_{K}]^{\mathsf{T}}$ are the Dirichlet parameters. These parameters govern how evenly or sparsely distributed the resulting distributions are. In particular, $\alpha\rightarrow 0$ favors distributions with nearly all mass concentrated on one of their components (i.e., sparse), $\alpha\rightarrow\infty$ favors near-uniform distributions, while for $\alpha=1$, all distributions are equally likely. This property is crucial in designing an algorithm which can automatically detect the model order of the PMF tensor (see Section III-D for more details).
@@
$p(\bm{a}_{n,r})$ $=\mathrm{Dir}(\bm{a}_{n,r}\,|\,\bm{\alpha}_{n,r})=C(\bm{\alpha}_{n,r})\prod_{i_{n}=1}^{I_{n}}a_{n,r,i_{n}}^{\alpha_{n,r,i_{n}}-1},$ (12)
where $\bm{\alpha}_{\lambda}=[\alpha_{\lambda,1},\dotsc,\alpha_{\lambda,R}]^{\mathsf{T}}$ and $\bm{\alpha}_{n,r}=[\alpha_{n,r,1},\dotsc,\alpha_{n,r,I_{n}}]^{\mathsf{T}}$ are the respective parameters for the prior distributions. In the absence of any prior information favoring one element over another, we choose symmetric Dirichlet distributions as priors, i.e., $\alpha_{\lambda,r}=\alpha_{\lambda}{\color[rgb]{0,0,0},}\,\,\forall r$, and $\alpha_{n,r,i_{n}}=\alpha_{n,r}{\color[rgb]{0,0,0},}\,\,\forall i_{n}$.
@@
Figure 1: Probabilistic graphical model representation of the joint distribution $p(\bm{Y},\bm{\Theta})$.
@@
×$\qquad\times p(\bm{\lambda})\cdot\prod_{n=1}^{N}\prod_{r=1}^{R}p(\bm{a}_{n,r}).$
[…]
The logarithm of the joint distribution is given by (cf. (5), (6), (11), and (12))
@@
$+(\alpha_{\lambda}-1)\sum_{r=1}^{R}\log\lambda_{r}$
$+\sum_{n=1}^{N}\sum_{r=1}^{R}\sum_{i_{n}=1}^{I_{n}}(\alpha_{n,r}-1)\log a_{n,r,i_{n}}+\mathrm{const},$
where ‘const’ refers to terms which do not depend on $\bm{Y}$ or $\bm{\Theta}$ and $a_{n,r,y_{n,t}}\triangleq\bm{A}_{n}(y_{n,t},r)$.
@@
Exact inference requires integration over all variables in $\bm{\Theta}$ in order to calculate the denominator. In the present case, this is a high-dimensional problem involving integration over the model parameters and summation over all possible (discrete) states of the latent variables, which is computationally infeasible. In the following, we resort to approximate Bayesian inference which yields closed-form expressions for the posterior distribution.
@@ III-B Posterior Approximation using Variational Inference
In variational inference (VI) [20], we seek a variational distribution $q(\bm{\Theta})$ that is closest to the true posterior distribution $p(\bm{\Theta}\,|\,\bm{Y})$ such that the Kullback-Leibler divergence (KLD) between them is minimized. The KLD is given by
@@
$=\log p(\bm{Y})-\underbrace{\int_{\bm{\Theta}}\bm{q}(\bm{\Theta})\log\frac{p(\bm{Y},\bm{\Theta})}{q(\bm{\Theta})}\mathrm{d}\bm{\Theta}}_{\mathcal{L}(q)}.$
A closer examination of (16) reveals that since the KLD is nonnegative and the logarithm of the marginal likelihood [^3] The marginal likelihood is defined as $p(\bm{Y})=\int_{\bm{\Theta}}p(\bm{Y},\bm{\Theta})\mathrm{d}\bm{\Theta}$. $\log p(\bm{Y})$ is independent of $\bm{\Theta}$, then maximizing the term $\mathcal{L}(q)$ is equivalent to minimizing $D(q(\bm{\Theta})\,\|\,p(\bm{\Theta}\,|\,\bm{Y}))$. Therefore, $\mathcal{L}(q)$ is a lower bound on $\log p(\bm{Y})$ and is usually referred to as the evidence lower bound (ELBO).
If all possible variational distributions $q(\bm{\Theta})$ are considered, then the KLD vanishes when $q(\bm{\Theta})=p(\bm{\Theta}\,|\,\bm{Y})$. However, since the true posterior distribution is intractable, some restrictions need to be imposed on $q(\bm{\Theta})$ to ensure tractability. A common way to restrict the family of variational distributions is to assume that $q(\bm{\Theta})$ can be factorized such that $q(\bm{\Theta})=\prod_{i}q_{i}(\bm{\Theta}_{i})$, where $\cup_{i}\bm{\Theta}_{i}=\bm{\Theta}$ and $\cap_{i}\bm{\Theta}_{i}=\varnothing$. In other words, the unknown quantities in $\bm{\Theta}$ are assumed to be mutually independent with each quantity being governed by a distinct variational distribution. This approach is called mean-field approximation. In our case, given that $\bm{\Theta}=\{\bm{Z},\bm{\lambda},\bm{A}_{1},\dotsc,\bm{A}_{N}\}$, the mean-field approximation yields
@@
Under this setup, it can be shown that the optimal distribution $q_{i}^{\ast}(\bm{\Theta}_{i})$ for the $i$-th component of $q(\bm{\Theta})$ (i.e., the distribution for which $\mathcal{L}(q)$ is largest) is
$\log q_{i}^{\ast}(\bm{\Theta}_{i})=\mathbb{E}_{q(\bm{\Theta}\backslash\bm{\Theta}_{i})}[\log p(\bm{Y},\bm{\Theta})]+\mathrm{const}.$ (18)
Here, $\mathbb{E}_{q(\bm{\Theta}\backslash\bm{\Theta}_{i})}[\cdot]$ denotes the expectation with respect to ${q(\bm{\bm{\Theta}})}$ over all components except $\bm{\Theta}_{i}$, while $\log p(\bm{Y},\bm{\Theta})$ is given in (14). The constant is found by normalizing the distribution $q_{i}^{\ast}(\bm{\Theta}_{i})$, i.e.,
@@
However, in most cases, the normalization constant will be readily apparent upon inspecting the resulting optimal distributions.
It can be seen from (18) that computing a particular variational distribution requires the knowledge of the other variational distributions. Therefore, the variational distributions are updated iteratively using a coordinate ascent algorithm (see Section III-D). In the following, we present the derivations of each variational distribution in (17).

### v2

@@ III Variational Bayesian PMF Estimation
In previous work on this topic, estimates of the CPD components $\bm{\lambda}$ and $\{\bm{A}_{n}\}_{n=1}^{N}$ (and, therefore, of the joint PMF tensor $\bm{\mathcal{X}}$) have been obtained from $\bm{Y}$ by first obtaining empirical estimates of lower-order marginal PMF tensors from $\bm{Y}$, followed by a coupled tensor factorization (e.g., [7], [9]). Alternatively, a maximum likelihood (ML) estimate of the PMF tensor was obtained by fitting the observed data to the CPD model (e.g., [12], [13]). While these approaches work quite well, they assume that the CPD rank $R$ is known, a scenario which rarely arises in practice. We therefore seek to formulate the problem within the Bayesian paradigm and design an algorithm to estimate the CPD components while automatically detecting the rank of $\bm{\mathcal{X}}$.
@@ III-A Model Specification
The Bayesian model is specified by first identifying the model parameters that one wishes to estimate, finding an expression for the likelihood of the observations given the model parameters, and then assigning prior distributions to the model parameters.
@@
$p\big(\bm{Y}\,\big|\,\bm{\lambda},\{\bm{A}_{n}\}_{n=1}^{N}\big)=\prod_{t=1}^{T}{\color[rgb]{0,0,0}\Bigg(}\sum_{r=1}^{R}\lambda_{r}\prod_{n=1}^{N}\bm{A}_{n}(y_{n,t},r){\color[rgb]{0,0,0}\Bigg)}.$ (4)
The naïve Bayes model structure allows us to introduce a latent variable $\bm{z}\in\mathbb{R}^{R}$ which contains the hidden mechanism that governs the observation $\bm{y}$. Let $\bm{z}=[z_{1},\dotsc,z_{R}]^{\mathsf{T}}$ where a particular element $z_{r}$ is equal to 1 and all other elements are equal to 0. Therefore, the elements of $\bm{z}$ satisfy $z_{r}\in\{0,1\}$ and $\sum_{r=1}^{R}z_{r}=1$. Each observation $\bm{y}$ is associated with a latent variable $\bm{z}$ which identifies the latent component from which $\bm{y}$ (and, therefore, $\bm{x}$) was drawn. In other words, if $\bm{y}$ is drawn from component $s\in\{1,\dotsc,R\}$ then $z_{r}=1$ if $r=s$ and $z_{r}=0$ if $r\neq s$. Defining $\bm{Z}=[\bm{z}_{1},\dotsc,\bm{z}_{T}]$ where $\bm{z}_{t}=[z_{1,t},\dotsc,z_{r,t}]^{\mathsf{T}}$, we can express the conditional distribution of $\bm{Y}$ given $\bm{Z}$ and $\{\bm{A}_{n}\}_{n=1}^{N}$ (up to a constant) as
@@
Next, we specify prior distributions for the component parameters $\bm{\lambda}$ and $\{\bm{A}_{n}\}_{n=1}^{N}$, which are subject to the probability simplex constraints (see Section II). An appropriate distribution for variables defined on a probability simplex is the Dirichlet distribution, which is a multivariate distribution over a set of random variables that are subject to nonnegativity and sum-to-one constraints. In particular, given $K$ random variables $\bm{\mu}=[\mu_{1},\dotsc,\mu_{K}]^{\mathsf{T}}$ , the Dirichlet distribution is given by (e.g., [3])
$\mathrm{Dir}(\bm{\mu}\,|\,\bm{\alpha})=C(\bm{\alpha})\prod_{k=1}^{K}\mu_{k}^{\alpha_{k}-1},~~{\color[rgb]{0,0,0}\bm{\mu}\in\Delta^{K-1}},$ (7)
where $\Delta^{K-1}\triangleq\big\{\bm{\mu}\in\mathbb{R}^{K}:\mu_{k}\geq 0,\sum_{k=1}^{K}\mu_{k}=1\big\}$ is the (K1)(K-1)-dimensional probability simplex. Furthermore, $C(\bm{\alpha})$ is the normalization constant, defined in terms of the standard gamma function [^3] For any real number $x$, the standard gamma function is defined as $\Gamma(x)=\int_{0}^{\infty}t^{x-1}\mathrm{e}^{-t}\mathrm{d}t$. $\Gamma(\cdot)$ as
@@
and $\bm{\alpha}=[\alpha_{1},\dotsc,\alpha_{K}]^{\mathsf{T}}$ are the Dirichlet parameters , subject to the constraint $\alpha_{k}>0,\forall k$. These parameters govern how evenly or sparsely distributed the resulting distributions are. In particular, $\alpha\rightarrow 0$ favors distributions with nearly all mass concentrated on one of their components (i.e., sparse), $\alpha\rightarrow\infty$ favors near-uniform distributions, while for $\alpha=1$, all distributions are equally likely (e.g., [3]). This property is crucial in designing an algorithm which can automatically detect the model order of the PMF tensor (see Section III-D for more details).
@@
$p(\bm{a}_{n,r})$ $=\mathrm{Dir}(\bm{a}_{n,r}\,|\,\bm{\alpha}_{n,r})=C(\bm{\alpha}_{n,r})\prod_{{\color[rgb]{0,0,0}i}=1}^{I_{n}}a_{n,r,{\color[rgb]{0,0,0}i}}^{\alpha_{n,r,{\color[rgb]{0,0,0}i}}-1},$ (12)
where $\bm{\alpha}_{\lambda}=[\alpha_{\lambda,1},\dotsc,\alpha_{\lambda,R}]^{\mathsf{T}}$ and $\bm{\alpha}_{n,r}=[\alpha_{n,r,1},\dotsc,\alpha_{n,r,I_{n}}]^{\mathsf{T}}$ are the respective parameters for the prior distributions. In the absence of any prior information favoring one element over another, we choose symmetric Dirichlet distributions as priors, i.e., $\alpha_{\lambda,r}=\alpha_{\lambda}{\color[rgb]{0,0,0},}\,\,\forall r$, and $\alpha_{n,r,{\color[rgb]{0,0,0}i}}=\alpha_{n,r}{\color[rgb]{0,0,0},}\,\,\forall{\color[rgb]{0,0,0}i}$.
@@
Figure 1: Probabilistic graphical model representation of the joint distribution $p(\bm{Y},\bm{\Theta})$. The local and global random variables are represented by the larger circles, whereas the hyperparameters are depicted by the smaller solid dots. The shaded circle denotes that $y_{t}$ is an observed realization. The plate represents $T$ instances, of which only $y_{t}$, $z_{t}$} are shown explicitly.
@@
$\qquad\times p(\bm{\lambda})\cdot\prod_{n=1}^{N}\prod_{r=1}^{R}p(\bm{a}_{n,r}).$
The probabilistic graphical model representation of (13) is shown in Fig. 1. The latent variables $\{\bm{z}_{t}\}_{t=1}^{T}$ are said to be local parameters since $\bm{z}_{t}$ only governs the data in the $t$-th context. On the other hand, $\bm{\lambda}$ and $\bm{a}_{n,r},\forall n,r$, are considered global parameters because their total number is fixed and does not scale with the dataset [22]. From (13), the logarithm of the joint distribution is given by (cf. (5), (6), (11), and (12))
@@
$\hskip-20.00003pt+(\alpha_{\lambda}-1)\sum_{r=1}^{R}\log\lambda_{r}$
$\hskip-20.00003pt+\sum_{n=1}^{N}\sum_{r=1}^{R}\sum_{{\color[rgb]{0,0,0}i}=1}^{I_{n}}(\alpha_{n,r}-1)\log a_{n,r,{\color[rgb]{0,0,0}i}}+\mathrm{const},$
where ‘const’ refers to terms which do not depend on $\bm{Y}$ or $\bm{\Theta}$, and $a_{n,r,y_{n,t}}\triangleq\bm{A}_{n}(y_{n,t},r)$.
@@
Exact inference requires integration over all variables in $\bm{\Theta}$ to calculate the denominator. In the present case, this is a high-dimensional problem that involves integration over the model parameters and summation over all possible (discrete) states of the latent variables, which is computationally infeasible. In the following, we resort to approximate Bayesian inference, which yields closed-form expressions for the posterior distribution.
@@ III-B Posterior Approximation using Variational Inference
TABLE I: Summary of priors, posteriors, update equations, and point estimates for variational Bayesian PMF estimation.
Parameter Prior Posterior Update Equations Point Estimate
$\bm{\lambda}\in\mathbb{R}^{R}$ $\mathrm{Dir}(\bm{\lambda}\,|\,\bm{\alpha}_{\lambda})$ $\mathrm{Dir}(\bm{\lambda}\,|\,\widetilde{\bm{\alpha}}_{\lambda})$ $\widetilde{\alpha}_{\lambda,r}=\alpha_{\lambda}+\sum_{t=1}^{T}\rho_{r,t}$ $\widehat{\lambda}_{r}=\frac{\widetilde{\alpha}_{\lambda,r}}{\sum_{j=1}^{R}\widetilde{\alpha}_{\lambda,j}}$
$\bm{a}_{n,r}\in\mathbb{R}^{I_{n}}$ $\mathrm{Dir}(\bm{a}_{n,r}\,|\,\bm{\alpha}_{n,r})$ $\mathrm{Dir}(\bm{a}_{n,r}\,|\,\widetilde{\bm{\alpha}}_{n,r})$ $\widetilde{\alpha}_{n,r,{\color[rgb]{0,0,0}i}}=\alpha_{n,r,{\color[rgb]{0,0,0}i}}+\sum_{t\in\Omega_{n,i_{n}}}\rho_{r,t}$ $\widehat{a}_{n,r,{\color[rgb]{0,0,0}i}}=\frac{\widetilde{\alpha}_{n,r,{\color[rgb]{0,0,0}i}}}{\sum_{j=1}^{I_{n}}\widetilde{\alpha}_{n,r,j}}$
$\bm{Z}\in\mathbb{R}^{R\times T}$ $\prod_{t=1}^{T}\prod_{r=1}^{R}\lambda_{r}^{z_{r,t}}$ $\prod_{t=1}^{T}\prod_{r=1}^{R}\rho_{r,t}^{z_{r,t}}$
$\gamma_{r,t}=\exp{\left\{\log\widetilde{\lambda}_{r}+\sum_{n=1}^{N}\log\widetilde{a}_{n,r,y_{n,t}}\right\}}$
$\rho_{r,t}=\frac{\gamma_{r,t}}{\sum_{j=1}^{R}\gamma_{j,t}}$
$\widehat{z}_{r,t}=\mathbb{E}[z_{r,t}]=\rho_{r,t}$
In variational inference (VI) [33, 22], we seek a variational distribution $q(\bm{\Theta})$ that is closest to the true posterior distribution $p(\bm{\Theta}\,|\,\bm{Y})$ such that the Kullback-Leibler divergence (KLD) between them is minimized. The KLD is given by
@@
$\hskip-50.00008pt=\log p(\bm{Y})-\underbrace{\int_{\bm{\Theta}}\bm{q}(\bm{\Theta})\log\frac{p(\bm{Y},\bm{\Theta})}{q(\bm{\Theta})}\mathrm{d}\bm{\Theta}}_{\mathcal{L}(q)}.$
A closer examination of (16) reveals that since the KLD is nonnegative and the marginal likelihood $p(\bm{Y})=\int_{\bm{\Theta}}p(\bm{Y},\bm{\Theta})\mathrm{d}\bm{\Theta}$ is independent of $\bm{\Theta}$, then maximizing the term $\mathcal{L}(q)$ is equivalent to minimizing $D(q(\bm{\Theta})\,\|\,p(\bm{\Theta}\,|\,\bm{Y}))$. Therefore, $\mathcal{L}(q)$ is a lower bound on $\log p(\bm{Y})$ and is usually referred to as the evidence lower bound (ELBO).
If all possible variational distributions $q(\bm{\Theta})$ are considered, then the KLD vanishes when $q(\bm{\Theta})=p(\bm{\Theta}\,|\,\bm{Y})$. However, since the true posterior distribution is intractable, some restrictions need to be imposed on $q(\bm{\Theta})$ to ensure tractability. A common way to restrict the family of variational distributions is to assume that $q(\bm{\Theta})$ can be factorized such that $q(\bm{\Theta})=\prod_{i}q_{i}(\bm{\Theta}_{i})$, where $\cup_{i}\bm{\Theta}_{i}=\bm{\Theta}$ and $\cap_{i}\bm{\Theta}_{i}=\varnothing$. Therefore, each of the unknown quantities $\bm{\Theta}_{i}$ is assumed to be governed by a distinct variational distribution $q_{i}(\bm{\Theta}_{i})$. This approach is called mean-field approximation [33, 22]. In our case, given that $\bm{\Theta}=\{\bm{Z},\bm{\lambda},\bm{A}_{1},\dotsc,\bm{A}_{N}\}$, the mean-field approximation yields
@@
Under the mean-field approximation, the optimized form of the $i$th factor $q_{i}^{\ast}(\bm{\Theta}_{i})$ is [3]
$\log q_{i}^{\ast}(\bm{\Theta}_{i})=\mathbb{E}_{q(\bm{\Theta}\backslash\bm{\Theta}_{i})}[\log p(\bm{Y},\bm{\Theta})]+\mathrm{const}{\color[rgb]{0,0,0},}$ (18)
which is derived by replacing $q(\bm{\Theta})$ in the expression for $\mathcal{L}(q)$ in (16) with its factorized form, fixing all but the $i$th factor, and maximizing $\mathcal{L}(q)$ with respect to $q_{i}(\bm{\Theta}_{i})$. Here, $\mathbb{E}_{q(\bm{\Theta}\backslash\bm{\Theta}_{i})}[\cdot]$ denotes the expectation with respect to $q(\bm{\bm{\Theta}})$} over all components except $\bm{\Theta}_{i}$, while $\log p(\bm{Y},\bm{\Theta})$ is given in (14). The constant is found by normalizing the distribution $q_{i}^{\ast}(\bm{\Theta}_{i})$
@@
However, in most cases, the constant will be readily apparent upon inspection of the resulting optimal distributions.
It can be seen from (18) that computing a particular variational distribution requires knowledge of the other variational distributions. Therefore, the variational distributions are updated iteratively using a coordinate ascent algorithm (see Section III-D). Since the ELBO is convex with respect to each factor $q_{i}(\bm{\Theta}_{i})$, coordinate ascent is guaranteed to converge to a local optimum [3]. In the following, we present the derivations of each variational distribution in (17). To keep the notation uncluttered, we will write $\mathbb{E}[\bm{\Theta}_{i}]$ in place of $\mathbb{E}_{q_{i}^{\ast}(\bm{\Theta}_{i})}[\bm{\Theta}_{i}]$ (i.e., the expectation of the $i$th unknown with respect to its variational distribution) where necessary.